OpenStudy (anonymous):

Translate the following situation into an inequality. DO NOT SOLVE The sum of three integers is no more than 110. If one integer is nine less than twice the first integer and the third integer is one-fourth the first integer. What are the integers.

OpenStudy (anonymous):

x+2x-9+x/4 <= 110

myininaya (myininaya):

x+y+z<=110 let x be the first integer y=2x-9 z=1/4*x x+(2x-9)+1/4*x<=110

OpenStudy (anonymous):

\[ <= means \le \] They're correct

OpenStudy (anonymous):

thanks everyone